# InstructDET: Diversifying Referring Object Detection with Generalized Instructions

### v1

@@ 3.4 Multi-Objects Expression Generation
Our process to produce summarized expression is shown in Fig. 4. For each object, we first concatenate all its related expressions with commas. Through this concatenation, we can obtain this object expression from different perspectives (i.e., different properties). Then, we use the text encoder in BERT (Devlin et al., 2018) to map this concatenated expression to a semantic embedding space. As such, we obtain embeddings of concatenated expressions from all objects. Then, we cluster these embeddings into indeterminate number of clusters by using DBSCAN (Ester et al., 1996) method. We use LLaMA to generate text from each cluster center. The details of using LLaMA to mine object commonalities are in Sec. E. The generated text indicates the summarized expression we aim to produce for multiple objects. We also note that there will be many clusters containing one point, which means that object represented by this point has no commonality with other objects.
Post Processing. After generating expressions for single and multiple objects, we verify and remove the expressions that repeatedly pertain to the same object. Then, we utilize LLaMA to diversify these expressions while preserving their original intents. We observe that for different objects in one image, the expression for each object may be similar to others. These expressions are ambiguous since we can not refer to a unique object based on their referring. Nevertheless, we transfer these expressions to refer to multi-objects since they express a set of objects in one image. This transfer further augments multi-object referring expressions. Finally, we collect these remaining expressions after post processing as instructions. Together with corresponding object bbxs and images, we construct our InDET dataset by incorporating diversified object detection instructions encompassing user intentions.
@@ 5 Referring Object Detection
In this section, we illustrate our model design for ROD task. We notice that ROD shares little difference with visual grounding (VG). First, ROD produces uncertain number of object bbxs (i.e., 0, 1, or multiple) based on one input instruction, as shown in Fig. 1. Second, ROD supports abstract and summarized object descriptions (e.g., “all objects on the table”) that do not clearly refer to specific objects such as “bottle”, “orange”, and “knife”. As recent VG models (Zhang et al., 2022; Liu et al., 2023b) require a one-by-one verification between visual objects and expression words, they are not able to execute such instructions. Motivated by the difference, we set up a conventional framework from DETR-based VG models (Zhu et al., 2022; Yan et al., 2023). Fig. 6 shows an overview of our DROD model. We illustrate key steps as follows:
@@
Given an image with text instruction, we use visual and text encoders (Dosovitskiy et al., 2021; Devlin et al., 2018) to obtain their embeddings. Then, we use a bi-directional cross attention module to perform multi-modality embedding fusion. For the fused visual embedding, we sent it to the transformer encoder and decoder structure (Zhu et al., 2020) with $N$ learnable queries as position priors  (Meinhardt et al., 2022; Wang et al., 2021). Then, the decoder produces $N$ instance proposals for further selection. For the fused text embedding, we pass it through a global average pooling and MLP for text2visual embedding space mapping. Finally, we use cosine similarity to match proposals and mapped text embedding. During the training stage, we use confidence loss and localization loss via supervised learning. During the inference stage, we select proposals whose matching scores are above a predefined threshold, which allows our model to produce arbitrary number of bbxs for diversified instruction execution. A detailed model design and experimental configurations are in Sec. G.
@@ 6 Experiments
We evaluate the ROD performance on standard VG benchmarks (i.e., RefCOCO, RefCOCO+, and RefCOCOg) and our InDET dataset. As illustrated in Sec. 4, the images with marked objects of our InDET dataset are collected from existing datasets while the instructions are significantly enriched. We split the training and test set of InDET following RefCOCO/g/+ where the test set contains 6.5k images with an increased number of instructions to 315K. Moreover, these instructions are assigned to 6 groups according to our guideline. The performance on each group reflects how VG methods perform when processing different aspects of user instructions. The comparing methods in our experiments are from recent VG methods including MDETR (Kamath et al., 2021), Grounding-DINO (Liu et al., 2023b) and UNINEXT (Yan et al., 2023). Due to page limit, we show the evaluation results on our InDET, our InDET with shuffled expression, and standard benchmarks. Model training, and ablation studies on partial LLaVA finetuning and visual prompt selection are provided in Sec. G and Sec. H. We also provide visual comparison results of these methods in Sec. A.
@@
MDETR ResNet101 GoldG,RefC 86.75 89.58 81.41 79.52 84.09 70.62 81.64 80.89
@@
Figure 7: Visual comparison results on a RefCOCO image. The instruction is “person who is NOT holding a bat”.
Besides evaluating our InDET test set, we compare our DROD model with existing VG methods (Zhu et al., 2022; Li & Sigal, 2021) on the standard VG benchmarks RefCOCO/g/+. Table 3 shows the evaluation results. Overall, our DROD model achieves favorable performance on these datasets. This is because our DROD model utilizes InDET dataset where diversified instructions improve model generalizations. By using a conventional ROD model, we improve the VG performance from the data diversity perspective. Furthermore, we show visual comparisons in Fig. 7 where the image is from RefCOCO. Given the instruction that contains opposite description (‘not’), both Grounding-DINO and UNINEXT can not comprehend this meaning by detecting the person not referred to. In comparison, our DROD model comprehends the instruction meaning and executes ROD correctly. This success is because of sufficient data coverage in our InDET dataset where diversified instructions improve model generalizations.
@@ 7 Concluding Remarks
We aim to push ROD into practical usage from a data-centric perspective. On one hand, we notice that current REC expressions are insufficient to encompass user detection intentions. On the other hand, foundation models have shown promising generalizations to simulate manual understanding and description abilities. To this end, we develop InstructDET that leverages foundation models to produce human-like expressions in REC, which tends to incorporate common user intentions into ROD training. As a result, our DROD model achieves favorable performance compared to existing VG methods. Since our data-centric method is only able to generalize based on recognized objects in bbxs, we only produce expressions in a closed set. In the future, we can combine our method with open-set object detectors to fully explore in-the-wild images (e.g., Internet images) for comprehensive user expression generation. We expect our DROD model to generalize as much as existing foundation models, and thus take a huge step towards completely solving ROD task.
@@
  • In Sec. B and Sec. C, we describe the process of generating expressions from foundation models in both multi-modality and single-modality pathways, respectively.
@@
  • In Sec. H, we present ablation studies investigating LLaVA fine-tuning in multi-modality pathway and visual prompting selection.
@@ Appendix A Visual Comparison Results
Appendix B Expression Generation via MULTI-MODALITY PATHWAY
We show the text prompts we use below in Table 4. For the LLaVA input, We randomly select one text prompt together with one image marked with object bbx. The inputs and samples generated from LLaVA are shown in Fig. 10.
{tcolorbox}
[colback=gray!10,colframe=black,width=14cm, arc=3mm, auto outer arc, ]
  • “Describe the objects in the red box.”
  • “Take a look at this image and describe What’s in the red box.”
  • “Please provide a description of the object in the red box.”
  • “Could you describe the contents in the red box of the image for me?”
  • “Use one sentence to index the objects in the red box.”
  • “Output a sentence describing the objects in the red box, so that people can locate the objects without ambiguity through this sentence.”
  • “Look carefully at the objects in the red box and describe them in one sentence to distinguish them from other objects.”
@@ Appendix B Expression Generation via MULTI-MODALITY PATHWAY
Figure 10: Examples of text prompts, images with marked object bbxs, and LLaVA outputs.
Appendix C Expression Generation via SINGLE-MODALITY PATHWAY
We choose an image from Objects365 as an example to illustrate the instruction generation pipeline via our single-modality generation pathway. This pathway consists of inputs, two steps, in-context samples, and LLaMA outputs, which are illustrated one-by-one in the following:
@@ Appendix C Expression Generation via SINGLE-MODALITY PATHWAY
Our input is an example image with object bbx coordinates, which is shown in Fig. 11.
@@
[colback=gray!10,colframe=black,width=14cm, arc=3mm, auto outer arc, ] [Fire Truck]
(1) vehicle, emergency vehicle, fire engine, parked outside the fire station, an essential part of the fire station’s resources, essential part of the fire station’s resources
(2) lined up in a neat row, ready for use, object parked in the row with other fire trucks, object with ladders and equipment
[Street Lights]
(1) light fixtures, outdoor lighting, two lights visible, providing illumination, source of illumination
(2) objects providing illumination, objects in the parking lot, objects providing a clear representation of the overall setting, objects providing light for the parking lot
@@
{tcolorbox}
[colback=gray!10,colframe=black,width=14cm, arc=3mm, auto outer arc, ]
  • “Describe the following image in detail”
  • “Provide a detailed description of the given image”
  • “Give an elaborate explanation of the image you see”
  • “Share a comprehensive rundown of the presented image”
  • “Offer a thorough analysis of the image”
  • “Explain the various aspects of the image before you”,
  • “Clarify the contents of the displayed image with great detail”
  • “Characterize the image using a well-detailed description”
  • “Break down the elements of the image in a detailed manner”
  • “Walk through the important details of the image”
  • “Portray the image with a rich, descriptive narrative”
  • “Narrate the contents of the image with precision”
  • “Analyze the image in a comprehensive and detailed manner”
  • “Illustrate the image through a descriptive explanation”
  • “Examine the image closely and share its details”
  • “Write an exhaustive depiction of the given image”
@@
[colback=gray!10,colframe=black,width=14cm, arc=3mm, auto outer arc, ] Prompt for LLaVA
Provide a detailed description of the given image, including objects: Street lights, Fire Truck
Output from LLaVA
Output 1st The image displays a parking lot outside a fire station where several fire trucks are parked. In total, there are five fire trucks of varying sizes, all lined up neatly in the lot. Additionally, there are two street lights visible in the scene, providing illumination for the area. The fire trucks appear ready for use and are an bessential part of the fire station’s resources.
Output 2nd The image depicts a parking lot outside a fire station, where several fire trucks are parked in a neat row. There are a total of five fire trucks of varying sizes, all aligned and ready for use. In addition to the fire trucks, there are two street lights visible in the scene, providing illumination for the parking lot. The overall setting gives a clear representation of the organized and prepared nature of the fire station.
@@
[colback=gray!10,colframe=black,width=14cm, arc=3mm, auto outer arc, ]
task_description = f""""
# Task description prompt
You are an AI visual assistant that can analyze a single image.
User will give you several sentences, describing the same image you are observing. In addition, specific interested object locations within the image are given, along with detailed coordinates. These coordinates are in the form of bounding boxes, represented as (x1, y1, x2, y2) with floating numbers ranging from 0 to 1. These values correspond to the left top x, left top y, right bottom x, and right bottom y.
Using the provided caption and bounding box information, give descriptions about the visual content of each interested objects as if you are seeing the image, as an assistant:
(1) give descriptions about the object itself, including object types, object functions, object counts, object locations, object actions, etc.
(2) give descriptions about the object and other objects, including the relative positions between objects, the interaction between objects in the image, etc.
Descriptions should be a series of phrases, not whole sentence. Give descriptions for specific interested objects only, do not centered on other objects.
Again, give descriptions centered on specific objects only.
image2text = f"""
# Image description of LLaVA, the following two paragraphs are from Table 7
The image displays a parking lot outside a fire station where several fire trucks are parked. In total, there are five fire trucks of varying sizes, all lined up neatly in the lot. Additionally, there are two street lights visible in the scene, providing illumination for the area. The fire trucks appear ready for use and are an bessential part of the fire station’s resources.
The image depicts a parking lot outside a fire station, where several fire trucks are parked in a neat row. There are a total of five fire trucks of varying sizes, all aligned and ready for use. In addition to the fire trucks, there are two street lights visible in the scene, providing illumination for the parking lot. The overall setting gives a clear representation of the organized and prepared nature of the fire station.
Street lights: [0.0, 0.23, 0.03, 0.26]
Fire truck: [0.05, 0.6, 0.21, 0.76], [0.19, 0.58, 0.37, 0.77], [0.33, 0.55, 0.61, 0.77], [0.56, 0.57, 0.74, 0.77], [0.72, 0.57, 1.0, 0.76]
# Python code together with above text prompts are directly sent to LLaMA
messages = [{"role": system, "content": task_description}]
for in_context_example in in_context_examples:
    messages.append({"role": user, "content": in_context_example["content"]})
    messages.append({"role": assistant, "content": in_context_example["response"]}
messages.append({"role": user, "content": image2text})
@@
[colback=gray!10,colframe=black,width=14cm,arc=3mm, auto outer arc, ]
Table 9: An in-context example. We manually prepare three in-context examples and show one here. On the top block there are image captions and object bbx coordinates. Note that the displayed image is only for image caption reference and is not used in practice. The bottom block shows our expected expressions for each object. The expressions listed in (1) focus on the object attribute itself, and listed in (2) focus on relationship between current object and other objects in this image.
Example image2text
Two children, a girl and a boy are practicing their writing.
Two children sit on a small seesaw in the sand.
Two children sitting on a teeter totter.
2 kids playing on a seesaw.
Two kids sit on a seesaw.
2 kids/two kids/two children: [0.09, 0.59, 0.49, 0.94]
girl: [0.09, 0.59, 0.23, 0.92]
boy: [0.34, 0.62, 0.49, 0.94]
seesaw/small seesaw/teeter totter: [0.06, 0.83, 0.57, 0.93]
sand: [0.01, 0.61, 1.0, 1.0]
  Example Response
[2 kids/two kids/two children]
(1) persons, children, two kids, children learning, kids playing, kids sitting, children practicing their writing
(2) 2 kids playing on a seesaw, Two kids sitting on a seesaw, a girl and a boy, two person on the teeter totter, kids by the sea
[girl]
(1) girl, kid on the left of the image, girl sitting, girl playing
(2) kid playing with the boy, girl sitting on the small seesaw, girl playing in the sand, girl reading a book, a friend of the boy on the right, the taller kid
[boy]
(1) boy, kid on the right, boy sitting, boy playing, boy practicing his writing
(2) kid playing with the girl, boy sitting on a teeter totter, a friend of the girl on the left, boy playing in the sand
[seesaw/small seesaw/teeter totter]
(1) small seesaw, teeter totter, item to be played on, common facilities in parks and playground, game of two people
(2) seesaw in the sand, item the kids are sitting on, item the girl is sitting on, item the boy is playing on
[sand]
(1) common by the sea, the background of the scene
(2) item on which the seesaw is placed, item on which the kids are standing

### v5

@@ 3.4 Multi-Objects Expression Generation
Our process to produce summarized expression is shown in Fig. 4. For each object, we first concatenate all its related expressions with commas. Through this concatenation, we can obtain this object expression from different perspectives (i.e., different properties). Then, we use the text encoder in BERT (Devlin et al., 2018) to map this concatenated expression to a semantic embedding space. As such, we obtain embeddings of concatenated expressions from all objects. Then, we cluster these embeddings into indeterminate number of clusters by using DBSCAN (Ester et al., 1996) method. We use LLaMA to generate text for clusters with multiple objects. The details of using LLaMA to mine object commonalities are in Sec. E. The generated text indicates the summarized expression we aim to produce for multiple objects.
Post Processing. After generating expressions for single and multiple objects, we verify and remove the repeated expressions that pertain to the same object. Then, we utilize LLaMA to further diversify these generated expressions while preserving their original intents, i.e., we use LLaMA to do synonymous rewriting of generated expressions. The prompt we use for synonymous rewriting is provided in Sec. I. We observe that for different objects in one image, the expression for one object may be similar to that of others. These expressions are ambiguous since we can not refer to a unique object based on their referring. Nevertheless, we transfer these expressions to refer to multi-objects since they express a set of objects in one image. This transfer further augments multi-object referring expressions. Finally, we collect these remaining expressions after post processing as instructions. Together with corresponding object bbxs and images, we construct our InDET dataset by incorporating diversified object detection instructions encompassing user intentions.
@@ 5 Referring Object Detection
In this section, we illustrate our model design for ROD task. We notice that ROD shares little difference with visual grounding (VG). First, ROD produces uncertain number of object bbxs (i.e., 0, 1, or multiple) based on one input instruction, as shown in Fig. 1. Second, ROD supports abstract and summarized object descriptions (e.g., “all objects on the table”) that do not clearly refer to specific objects such as “bottle”, “orange”, and “knife”. As recent VG models (Zhang et al., 2022; Liu et al., 2023c) require a one-by-one verification between visual objects and expression words, they are not able to execute such instructions. Motivated by the difference, we set up a conventional framework from DETR-based VG models (Zhu et al., 2022; Yan et al., 2023). Fig. 6 shows an overview of our DROD model. We illustrate key steps as follows:
@@
Given an image with text instruction, we use visual and text encoders (Dosovitskiy et al., 2021; Devlin et al., 2018; Ge et al., 2023) to obtain their embeddings. Then, we use a bi-directional cross attention module to perform multi-modality embedding fusion. For the fused visual embedding, we sent it to the transformer encoder and decoder structure (Zhu et al., 2020) with $N$ learnable queries as position priors  (Meinhardt et al., 2022; Ge et al., 2021). Then, the decoder produces $N$ instance proposals for further selection. For the fused text embedding, we pass it through a global average pooling and MLP for text2visual embedding space mapping. Finally, we use cosine similarity to match proposals and mapped text embedding. During the training stage, we use confidence loss and localization loss via supervised learning. During the inference stage, we select proposals whose matching scores are above a predefined threshold, which allows our model to produce arbitrary number of bbxs for diversified instruction execution. More details are shown in Sec. G.
@@ 6 Experiments
We evaluate the ROD performance on standard VG benchmarks (i.e., RefCOCO, RefCOCO+, and RefCOCOg) and our InDET dataset. As illustrated in Sec. 4, the images with marked objects of our InDET dataset are collected from existing datasets while the instructions are significantly enriched. We split the training and test set of InDET following RefCOCO/g/+ where the test set contains 6.5k images with an increased number of instructions to 315K. Moreover, these instructions are assigned to 6 groups according to our guideline. The performance on each group reflects how VG methods perform when processing different aspects of user instructions. The comparing methods in our experiments are from recent VG methods including MDETR (Kamath et al., 2021), Grounding-DINO (Liu et al., 2023c) and UNINEXT (Yan et al., 2023). Due to page limit, we show the evaluation results on our InDET, our InDET with shuffled expression, and standard benchmarks. Model training, and ablation studies on partial LLaVA finetuning and visual prompt selection are provided in Sec. G and Sec. H. We also provide visual comparison results of these methods in Sec. A. A video demo showing the practical usage of our DROD model is in our webpage.
@@
MDETR ResNet101 GoldG,CC,RefC 86.75 89.58 81.41 79.52 84.09 70.62 81.64 80.89
@@
Besides evaluating our InDET test set, we compare our DROD model with existing VG methods (Zhu et al., 2022; Li & Sigal, 2021) on the standard VG benchmarks RefCOCO/g/+. Table 3 shows the evaluation results. Overall, our DROD model achieves favorable performance on these datasets. This is because our DROD model utilizes InDET dataset where diversified instructions improve model generalizations. By using a conventional ROD model, we improve the VG performance from the data diversity perspective.
Figure 7: Our InDET dataset improves logic reasoning of ROD models. In (a), existing models trained with our InDET dataset show superior results compared to other datasets. In (b), we show visual comparisons by using the same DROD model but with different training datasets.
In addition to the overall precision comparisons, we evaluate how our dataset improves logic reasoning and instruction comprehension of existing models. Specifically, we select 2k test samples from our InDET test dataset where logic reasoning on instructions is required for object detection. For each model (i.e., MDETR, G-DINO, or UNINEXT), we train it by using different datasets (i.e., RefCOCO, Flicker, or InDET) and show the performance comparison on our 2k test samples. Fig. 7(a) shows the evaluation results where each model trained with our InDET dataset outperforms the same model trained with other datasets. In Fig. 7(b), we show visual comparisons by using our DROD model but with different training sets. It shows that using original datasets, the model tends to ground keywords rather than preform multi-modal reasoning based on instructions. In comparison, by training with our InDET, the model well interprets instruction meaning and conduct logic reasoning across languages and visual images.
@@ 7 Concluding Remarks
We aim to push ROD into practical usage from a data-centric perspective. On one hand, we notice that current REC expressions are insufficient to encompass user detection intentions. On the other hand, foundation models have shown promising generalizations to simulate manual understanding and description abilities. To this end, we develop InstructDET that leverages foundation models to produce human-like expressions in REC, which tends to incorporate common user intentions into ROD training. As a result, our DROD model achieves favorable performance compared to existing VG methods. In the future, we can combine our method with open-set object detectors to fully explore in-the-wild images (e.g., Internet images) for comprehensive user expression generation. We expect our DROD model to generalize as much as existing foundation models, and thus take a huge step towards completely solving ROD task.
Acknowledgement
This paper is supported by the National Natural Science Foundation of China under Grants (62073245, 62173248,62233013). Shanghai Science and Technology Innovation Action Plan (22511104900). Shanghai Municipal Science and Technology Major Project (2021SHZDZX0100) and the Fundamental Research Funds for the Central Universities.
@@ Acknowledgement
  • In Sec. B and Sec. C, we describe the process of generating expressions from foundation models in both global prompt and local prompt pipelines, respectively.
@@
  • In Sec. H, we present ablation studies investigating LLaVA fine-tuning in local prompt pipeline and visual prompting selection.
  • In Sec. I, we provide the prompt we used in post processing.
  • In Sec. J, we add supplementary evaluation results to Table. 2 and Table. 3.
  • In Sec. K, we add the time consumption of main steps in our instruction generation procedure.
@@ Appendix A Visual Comparison Results
Appendix B Expression Generation via LOCAL PROMPT PIPELINE
B.1 Prompt and Generation Display
We show the text prompts we use below in Table 4. For the LLaVA input, We randomly select one text prompt together with one image marked with object bbox. The inputs and samples generated from LLaVA are shown in Fig. 10.
“Describe the objects in the red box.” “Take a look at this image and describe What’s in the red box.” “Please provide a description of the object in the red box.” “Could you describe the contents in the red box of the image for me?” “Use one sentence to index the objects in the red box.” “Output a sentence describing the objects in the red box, so that people can locate the objects without ambiguity through this sentence.” “Look carefully at the objects in the red box and describe them in one sentence to distinguish them from other objects.”
@@ B.1 Prompt and Generation Display
Figure 10: Examples of text prompts, images with marked object bboxs, and LLaVA outputs.
B.2 Fine-tuning Settings
Dataset Construction
We employed the entire dataset from RefCOCO/+/g for our fine-tuning endeavors, which primarily comprises triplets of images, bounding boxes, and instructions. The detection boxes are overlaid onto the images, serving as the visual input for LLAVA, while the instructions from the dataset are used as supervised outputs. The prompts are randomly selected from Table 4. The overall dataset consists of 386k instructions paired with 66k images. We chose not to utilize additional REC datasets for fine-tuning, as our fine-tuning objective is solely directed towards enabling LLAVA to comprehend our task and acquire the ability to describe targets within red detection boxes.
Training Process
Firstly, it is crucial to emphasize that in this paper, LLAVA does not refer to the model presented in the original LLAVA paper. Instead, it broadly denotes a multimodal paradigm where image information is abstracted into several tokens and inputted into a Large Language Model (LLM) for cross-modal attention with textual tokens. Our LLAVA architecture utilizes Q-Former to extract 32 tokens from the VIT features of the image, and these 32 tokens are subsequently mapped to the same feature space as textual tokens through a linear layer. During fine-tuning, only the final mapping linear layer is trained. This fine-tuning approach enables the model to learn the salience of the target object within the red bounding box in the input image while ensuring that the overall model avoids catastrophic forgetting. We initialize our model with the weights of MiniGPT-4.
Hyperparameter Configuration and Implementation Details
We employed Vicuna 13B as our Large Language Model (LLM). The batch size and number of epochs were set to 32 and 30, respectively. The initial learning rate was set to $3\times 10^{-5}$. Model training was conducted using 4 threads on a single A100 80GB GPU.
B.3 Chain-of-Thought and Reflection
After fine-tuning, the LLAVA model still exhibits a certain degree of hallucinations, manifesting as the generation of imaginary objects. Taking inspiration from the concept of thought chains in large language models, we manually constructed a thought chain specifically tailored for the task of describing target objects. In Table 5, we illustrate the prompts at each step of the thought chain: 1) Inquire about the category of the target object. 2) Inquire about the attributes possessed by the target object. 3) Inquire about the objects surrounding the target object. 4) Inquire about the relationship between the target object and its surrounding objects. Through these four steps, we can attain a comprehensive understanding of LLAVA’s perception of the target object. However, hallucinations may still persist in the thought chain; therefore, in step 5), we prompt the model to reexamine the image and correct any errors. Finally, in step 6), we prompt LLAVA to produce the ultimate description of the target object.
1) “What is the object inside the red bounding box?” 2) “What are the attributes of the object within the red bounding box?” 3) “Which objects are around the target object in the red box?” 4) “What is the relationship between the object inside the red bounding box and the surrounding objects?” 5) “Please review the image once again, and if there are any inaccuracies in your previous answers regarding the object’s attributes and relationships, kindly correct them.” 6) “Look carefully at the objects in the red box and describe them in one sentence to distinguish them from other objects.”
Table 5: Text prompts for guiding chain-of-thought and reflection.
Appendix C Expression Generation via GLOBAL PROMPT PIPELINE
We choose an image from Objects365 as an example to illustrate the instruction generation pipeline via our global prompt pipeline. This pipeline consists of inputs, two steps, in-context samples, and LLaMA outputs, which are illustrated one-by-one in the following:
@@ Appendix C Expression Generation via GLOBAL PROMPT PIPELINE
Our input is an example image with object bbox coordinates, which is shown in Fig. 11.
@@
[Fire Truck] (1) vehicle, emergency vehicle, fire engine, parked outside the fire station, an essential part of the fire station’s resources, essential part of the fire station’s resources (2) lined up in a neat row, ready for use, object parked in the row with other fire trucks, object with ladders and equipment [Street Lights] (1) light fixtures, outdoor lighting, two lights visible, providing illumination, source of illumination (2) objects providing illumination, objects in the parking lot, objects providing a clear representation of the overall setting, objects providing light for the parking lot
@@
Figure 11 and Table 6 depict the original input and final output of the global prompt pipeline, with the intermediary processing steps outlined below. Initially, the generation of the desired caption description for images lacking dense captions is required (Step 1). Subsequently, with the image fully textualized, we employ a detailed prompt design and seed example design to instruct LLAMA in outputting instructions describing the target object (Step 2).
@@
“Describe the following image in detail” “Provide a detailed description of the given image” “Give an elaborate explanation of the image you see” “Share a comprehensive rundown of the presented image” “Offer a thorough analysis of the image” “Explain the various aspects of the image before you”, “Clarify the contents of the displayed image with great detail” “Characterize the image using a well-detailed description” “Break down the elements of the image in a detailed manner” “Walk through the important details of the image” “Portray the image with a rich, descriptive narrative” “Narrate the contents of the image with precision” “Analyze the image in a comprehensive and detailed manner” “Illustrate the image through a descriptive explanation” “Examine the image closely and share its details” “Write an exhaustive depiction of the given image”
@@
Prompt for LLaVA Provide a detailed description of the given image, including objects: Street lights, Fire Truck Output from LLaVA Output 1st The image displays a parking lot outside a fire station where several fire trucks are parked. In total, there are five fire trucks of varying sizes, all lined up neatly in the lot. Additionally, there are two street lights visible in the scene, providing illumination for the area. The fire trucks appear ready for use and are an bessential part of the fire station’s resources. Output 2nd The image depicts a parking lot outside a fire station, where several fire trucks are parked in a neat row. There are a total of five fire trucks of varying sizes, all aligned and ready for use. In addition to the fire trucks, there are two street lights visible in the scene, providing illumination for the parking lot. The overall setting gives a clear representation of the organized and prepared nature of the fire station.
@@
task_description = f"""" # Task description prompt ## Establishing fundamental roles and task localization for LLM. You are an AI visual assistant that can analyze a single image. ## Elucidating the form and significance of input information. User will give you several sentences, describing the same image you are observing. In addition, specific interested object locations within the image are given, along with detailed coordinates. These coordinates are in the form of bounding boxes, represented as (x1, y1, x2, y2) with floating numbers ranging from 0 to 1. These values correspond to the left top x, left top y, right bottom x, and right bottom y. ## Explicating the output content and the encapsulated information. Using the provided caption and bounding box information, give descriptions about the visual content of each interested objects as if you are seeing the image, as an assistant: (1) give descriptions about the object itself, including object types, object functions, object counts, object locations, object actions, etc. (2) give descriptions about the object and other objects, including the relative positions between objects, the interaction between objects in the image, etc. ## Emphasizing common output issues. Descriptions should be a series of phrases, not whole sentence. Give descriptions for specific interested objects only, do not centered on other objects. Again, give descriptions centered on specific objects only. """ image2text = f""" # Image description of LLaVA, the following two paragraphs are from Table 8 The image displays a parking lot outside a fire station where several fire trucks are parked. In total, there are five fire trucks of varying sizes, all lined up neatly in the lot. Additionally, there are two street lights visible in the scene, providing illumination for the area. The fire trucks appear ready for use and are an bessential part of the fire station’s resources. The image depicts a parking lot outside a fire station, where several fire trucks are parked in a neat row. There are a total of five fire trucks of varying sizes, all aligned and ready for use. In addition to the fire trucks, there are two street lights visible in the scene, providing illumination for the parking lot. The overall setting gives a clear representation of the organized and prepared nature of the fire station. Street lights: [0.0, 0.23, 0.03, 0.26] Fire truck: [0.05, 0.6, 0.21, 0.76], [0.19, 0.58, 0.37, 0.77], [0.33, 0.55, 0.61, 0.77], [0.56, 0.57, 0.74, 0.77], [0.72, 0.57, 1.0, 0.76] """" # Python code together with above text prompts are directly sent to LLaMA messages = [{"role": system, "content": task_description}] for in_context_example in in_context_examples:     messages.append({"role": user, "content": in_context_example["content"]})     messages.append({"role": assistant, "content": in_context_example["response"]} messages.append({"role": user, "content": image2text})
@@
[Uncaptioned image] Example image2text Two children, a girl and a boy are practicing their writing. Two children sit on a small seesaw in the sand. Two children sitting on a teeter totter. 2 kids playing on a seesaw. Two kids sit on a seesaw. 2 kids/two kids/two children: [0.09, 0.59, 0.49, 0.94] girl: [0.09, 0.59, 0.23, 0.92] boy: [0.34, 0.62, 0.49, 0.94] seesaw/small seesaw/teeter totter: [0.06, 0.83, 0.57, 0.93] sand: [0.01, 0.61, 1.0, 1.0]   Example Response [2 kids/two kids/two children] (1) persons, children, two kids, children learning, kids playing, kids sitting, children practicing their writing (2) 2 kids playing on a seesaw, Two kids sitting on a seesaw, a girl and a boy, two person on the teeter totter, kids by the sea [girl] (1) girl, kid on the left of the image, girl sitting, girl playing (2) kid playing with the boy, girl sitting on the small seesaw, girl playing in the sand, girl reading a book, a friend of the boy on the right, the taller kid [boy] (1) boy, kid on the right, boy sitting, boy playing, boy practicing his writing (2) kid playing with the girl, boy sitting on a teeter totter, a friend of the girl on the left, boy playing in the sand [seesaw/small seesaw/teeter totter] (1) small seesaw, teeter totter, item to be played on, common facilities in parks and playground, game of two people (2) seesaw in the sand, item the kids are sitting on, item the girl is sitting on, item the boy is playing on [sand] (1) common by the sea, the background of the scene (2) item on which the seesaw is placed, item on which the kids are standing
Table 10: An in-context example. We manually prepare three in-context examples and show one here. On the top block there are image captions and object bbox coordinates. Note that the displayed image is only for image caption reference and is not used in practice. The bottom block shows our expected expressions for each object. The expressions listed in (1) focus on the object attribute itself, and listed in (2) focus on relationship between current object and other objects in this image.
The in-context example in Table 10 follows the task description in Table 9 and primarily serves to standardize the output format of LLAMA. The desired output format is intended to resemble the Example Response presented in Table 9. [name] denotes the target currently being described, where the descriptions in (1) are relatively simple, pertaining only to the inherent attributes, and those in (2) entail more complex descriptions involving surrounding objects. This facilitates accurate correspondence between each description and the object it pertains to when parsing the output descriptions.
It is noteworthy that the prompt design is manually crafted. It draws significant inspiration from the prompt framework used by LLAVA, with specific modifications and supplements tailored to the requirements of our task. The entire prompt design underwent numerous iterations based on the output results, ensuring the stability of the output format and style.